# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversational Agents

### v4

@@ 4 DialSim
Building on LongDialQA, our simulator features an agent taking on the role of a main character in a dialogue (i.e., Ross, Sheldon, and Michael). Throughout the simulation, an agent is randomly asked questions by other characters that must be answered accurately within a time limit (§ 4.2).
4.1 Definition
Let the $k$-th utterance of the $n$-th session be denoted as $u_{n,k}$, and the $n$-th session consisting of $r$ utterances be $\mathcal{S}_{n}=\{\{u_{n,i}\}_{i=1}^{r},d_{n}\}$, where $d_{n}$ is the date of $\mathcal{S}_{n}$. The sub-session including up to the $k$-th utterance of the $n$-th session is $\mathcal{S}_{n,k}=\{\{u_{n,i}\}_{i=1}^{k},d_{n}\}$. The entire dialogue consisting of $N$ sessions is denoted as $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{N}$. The agent’s memory up to the $k$-th utterance of the $n$-th session is $\mathcal{M}_{n,k}$. The agent answering question $q_{n,m,c}$ asked by character $c$ in the $m$-th utterance of the $n$-th session using the memory is $a_{n,m}=Agent(\mathcal{M}_{n,m},q_{n,m,c})$.
Input: $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{N}$, Time interval $t$, Agent
@@ 15                   $\mathcal{M}_{n,k}\leftarrow UpdateMemory(\mathcal{M}_{n,k-1},$ $u_{n,k},d_{n},t)$;
17                         $a_{n,m}\leftarrow AgentAnswer(\mathcal{M}_{n,m},$ $q_{n,m,c},d_{n},t)$;
@@ 4.2 Simulator
Algorithm 1 outlines the simulation process of DialSim, designed to emulate a real-time conversation. In this simulator, each participant’s utterance (including the agent’s) occurs at a predefined time interval (same as time limit), and the agent should update its memory within this interval. [^5] The memory can be incrementally updated in various ways (e.g., by storing each utterance separately or by summarizing the session up to the current utterance). A detailed discussion of these methods is provided in § 5.2. If updating the memory is not completed within the interval, the simulator will move on to the next utterance (Line 13). During the simulation, other characters ask questions (selected from LongDialQA) to the agent (Line 8-10), except in sessions where the agent is the only one talking (Line 5-6). The timing to ask a question is chosen randomly within the session (Line 8), and the speaker who asks the question is also chosen randomly. However, to make the simulation realistic, it is crucial to ensure that the chosen speaker is still present and hasn’t left the session. We achieved this by randomly choosing from characters who were present within three turns of the agent’s last utterance (Line 9). Then, a question is randomly selected and asked in the style of the corresponding speaker (Line 10). The agent then must respond to the question using its memory, all within the time limit (Line 15). The prompt for the response is created by combining the question with the dialogue history stored in the memory. If the response is not completed within the time limit, it will be considered a failure, and the simulator will move on to the next utterance. The prompt we used is provided in Appendix F.
@@ 5 Experiments
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 38.53 (0.89) 32.65 (2.65) 49.04 (1.67) 40.27 (1.36) 40.10 (0.75) 44.36 (2.36) 42.53 (1.26) ChatGPT-3.5 31.82 (1.31) 25.58 (1.78) 39.70 (1.86) 32.09 (0.84) 32.06 (1.60) 36.84 (1.77) 36.69 (1.25) Gemini 1.0 pro 2.96 (0.31) 28.77 (1.83) 25.07 (2.40) 35.27 (1.80) 34.22 (0.49) 31.83 (0.41) 35.75 (2.93) Open Tülu 2-70B 0.37 (0.15) 20.94 (0.75) 20.27 (0.99) 19.75 (0.08) 31.76 (1.84) 10.15 (0.55) 18.87 (0.30) Tülu 2-7B 0.84 (0.15) 12.68 (0.24) 19.58 (1.04) 26.84 (0.85) 14.08 (0.89) 17.39 (1.37) 25.21 (1.28) Llama3.1-70B 0.60 (0.06) 31.08 (1.21) 0.55 (0.12) 16.26 (5.05) 39.00 (0.30) 2.26 (0.42) 20.14 (0.22) Llama3.1-8B 28.82 (1.94) 27.12 (0.95) 34.14 (0.85) 30.91 (0.63) 29.76 (1.31) 33.25 (0.57) 24.48 (0.60) Mixtral-8x7B 1.88 (0.26) 16.84 (0.95) 26.23 (0.90) 17.11 (1.94) 17.94 (1.32) 26.78 (1.04) 15.40 (1.39) Mistral-7B 2.82 (0.46) 24.22 (2.04) 33.07 (1.01) 29.29 (1.76) 28.30 (1.93) 29.15 (1.67) 25.41 (1.53) Gemma-7B 16.60 (0.84) 22.11 (1.73) 24.30 (2.04) 18.33 (1.37) 26.42 (2.48) 22.54 (0.78) 18.80 (0.64) Gemma-2B 0.68 (0.20) 24.06 (2.03) 24.22 (1.34) 25.79 (1.00) 25.31 (1.55) 24.48 (1.62) 25.78 (1.12)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@ 5.1 Experimental Setting
To efficiently and accurately evaluate the agents’ dialogue understanding abilities, we used a multiple-choice format for the questions in the experiments. Table 1 shows the statistics for LongDialQA, revealing a notable difference between the number of answerable and unanswerable questions. To ensure a balanced distribution of correct answers during the simulation, 20% of the questions were intentionally designed to be unanswerable, with each question offering five possible choices. In addition to the multiple-choice format, we also offer an option to use an open-ended format, allowing users to choose their preferred question format.
@@ 5.2 Baselines
We experimented with two methods for using an agent’s memory. The first method, namely Base LLM, is to simply prefix latest utterances as much allowed by the model’s context length. The second method, namely RAG-based, employs a retriever to search for relevant dialogue history from the agent’s memory (external storage) and includes it in the prompt (Lewis et al., 2020). This method can be broken down into three ways for storing dialogue history: each speaker’s utterance individually, the entire session, and a summarized version of each session (denoted as Utterance, Session Entire, and Session Sum. in Table 2). The retrieval from the memory was performed using BM25 (Robertson et al., 2009) and cosine similarity with the OpenAI embeddings (OpenAI, 2024c).
For the agents to be tested, we used both API-based models (i.e., Gemini-1.0 Pro, 1.5 Pro (Team et al., 2023; Reid et al., 2024), Claude 3 Opus (Anthropic, 2024), ChatGPT-3.5, 4o, 4o-mini (OpenAI, 2023b, 2024b, 2024a)) and open-source models (i.e., Tülu 2-7B, 70B (Ivison et al., 2023), Llama3.1-8B, 70B (Meta, 2024), Mistral-7B, 8x7B (Jiang et al., 2023, 2024), and Gemma-2B, 7B (Team et al., 2024)). [^6] Gemini-1.5 Pro, Claude 3 Opus, and ChatGPT-4o were evaluated only in the BM25-Session Entire and oracle setting to measure their performance upper bound due to their high prices. The experimental results can be found in Appendix K. To emulate conversational settings, we used chat templates for instruction-tuned models or directly used chat models.
@@ 5.3 Results
Overall Performance   Table 2 shows that API-based models outperformed open-source models due to their superior inference capabilities and faster response times in our setting. However, the performances of all baselines were below 50%, suggesting that current LLMs have limitations in their ability to serve as conversational agents for long-term multi-party dialogues. The experimental results for Friends, The Big Bang Theory, and The Office exhibited similar trends. The detailed results are described in Appendix H.
Model Base LLM BM25 OpenAI Embedding UT SE SS UT SE SS ChatGPT-4o-mini 38.91 34.44 49.21 42.23 38.91 43.64 42.40 ChatGPT-3.5 31.81 26.91 39.45 32.77 32.41 35.78 35.98 Gemini 1.0 pro 28.36 28.10 39.90 34.11 34.26 30.93 33.96 Llama3.1-70B 36.36 31.84 43.17 43.81 39.85 43.17 48.49 Llama3.1-8B 28.78 29.89 34.70 33.93 31.63 32.91 35.59 Mixtral-8x7B 42.19 31.84 46.47 32.31 35.51 41.24 34.18 Mistral-7B 32.93 28.20 35.09 30.16 30.12 31.00 30.10 $\dagger$: Limited the maximum context length to 8k tokens.
@@
Storing the entire session consistently outperforms other history storing methods, as shown in Table 3. This is because individual utterances lack adequate context, and crucial information may be lost during summarization. However, Llama3.1 models achieved the best performance when using Session Sum. as a history saving method, owing to their strong summarization capabilities. Additionally, contrary to our expectations, Mixtral’s Base LLM (i.e., without history retrieval) outperforms some retrieval-based models in settings with unlimited time. This is due to Mixtral’s context length of 32k tokens, which is long enough to accommodate half a season of the script, allowing it to utilize a longer dialogue history than some of the other baselines. However, in a setting with a time limit, Mixtral’s performance significantly drops due to its long inference time. Therefore, for a conversational agent to converse in real-time, it is necessary to select a reasonably appropriate length of dialogue history.
Advanced techniques for storing and retrieving history are essential to engage in long-term multi-party dialogues. We conducted experiments under the oracle setting, where agents were given evidence sessions along with their dates (see Figure 2). Under these conditions, Llama3.1-70B achieved a top performance of 69.86% in an unlimited time scenario, outperforming the best RAG-based method by 21.37%. This significant performance gap highlights the importance of effective memory management techniques. Detailed experimental results are provided in Appendix K.
TKG-based questions present a greater challenge than fan quiz-based ones, with two-hop questions being particularly difficult. To assess the difficulty levels across different question types, we conducted an error analysis on ChatGPT-4o-mini, based on BM25-Session Entire, which showed the highest performance. The results showed that fan quiz-based questions had an accuracy of 58.80%, while TKG-based questions scored lower at 46.40%, highlighting the greater difficulty of TKG-based questions. Breaking down TKG-based questions further, one-hop questions had a performance of 66.67%, whereas two-hop questions had a performance of 13.53%, underscoring the challenge of two-hop questions. Furthermore, even in the oracle setting, while the performance of one-hop questions increased to 84.05%, two-hop questions remained at 28.45%. This indicates that two-hop questions are challenging not only in terms of history retrieval but also in reasoning across the given sessions.
Model Original Swapping Names New Names Llama3.1-70B 48.49 39.00 ($\downarrow$9.49) 44.27 ($\downarrow$4.22) Llama3.1-8B 35.59 31.59 ($\downarrow$4.00) 32.23 ($\downarrow$3.36) Mixtral-8x7B 46.47 37.72 ($\downarrow$8.75) 39.98 ($\downarrow$6.49) Mistral-7B 35.09 30.65 ($\downarrow$4.44) 34.59 ($\downarrow$0.50)
@@
Adversarial testing is necessary to accurately evaluate dialogue understanding in conversational agents. We conducted further experiments for the adversarial test by altering the names of the characters in two ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). The results shown in Table 4 indicated a significant drop in overall performance compared to the original setup. This decline is attributed to the agents relying not only on the dialogue history but also on their pre-trained knowledge when answering questions. Additionally, the performance decrease was more pronounced when names were swapped compared to when new names were assigned. This suggests that new names represent new information, while mixed names in the dialogue history conflicted with the pre-trained knowledge, leading to reduced reasoning ability. The detailed experimental results are provided in Appendix L.
@@ 6 Conclusion
In this paper, we introduce DialSim, a simulator designed to evaluate the capabilities of conversational agents in understanding long-term, multi-party dialogues in real-time settings. To run DialSim, we first constructed LongDialQA, a dataset based on dialogues from well-known TV show scripts. LongDialQA also includes questions derived from fan quizzes and a temporal knowledge graph, enabling a comprehensive assessment of conversational agents. Using DialSim, we evaluated the latest conversational agents and uncovered significant limitations in their ability to effectively handle complex, multi-party, long-term dialogues in real-time scenarios.
@@
Despite its strengths, our simulator has two main limitations. First, while the questions and answers are logically paired for accurate evaluation, the random selection of questions could introduce a bit of awkwardness during conversations. Second, while we considered incorporating industry-specific dialogues such as chat logs from customer service or retail, where conversational agents could be used for business purposes, these dialogue datasets are usually proprietary and not publicly accessible. In future developments, we will focus on enhancing the natural flow of interactions and creating simulators that are applicable to real-world industries.
@@ Appendix A LLM’s Prior Knowledge of the TV shows
We asked ChatGPT-4o to explain the plot of specific episodes of Friends. It accurately described the plots, as shown in Figure 45. Notably, it provided these answers without any web browsing, suggesting that ChatGPT-4o might have learned about these TV shows during its pre-training process.
@@ Appendix C Question Generation Based on Fan Quizzes
First, we collected quizzes for each season and episode of Friends, The Big Bang Theory, and The Office from the FunTrivia website. For each episode $p_{i}$ in $Script_{pre}$, we used ChatGPT-4 to determine if the crawled questions ${CrQ}_{i}=\{q_{i,0},q_{i,1},...,q_{i,l}\}$ could be answered using only $p_{i}$. If a question $q_{i,m}$ could be answered, ChatGPT-4 identified the scenes $ES_{i,m}$ that provide evidence for the answer, compiling them into $Q_{i}=\{(q_{i,m},ES_{i,m})\}_{m=0}^{l}$. Subsequently, the authors reviewed each $ES_{i,m}$, made necessary corrections, and annotated whether a single scene from $ES_{i,m}$ was sufficient to answer $q_{i,m}$ or if multiple scenes were needed to be considered simultaneously. For each $s_{i,k}$ within $p_{i}$, we assessed the answerability of the questions in $Q_{i}$.

### v5

@@ 4 DialSim
Building on LongDialQA, DialSim features a conversation system taking on the role of a main character in a dialogue (i.e., Ross, Sheldon, and Michael). Throughout the simulation, the system is asked questions by other characters that must be answered accurately within a time limit (§ 4.1).
Input: $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{N}$, Time interval $t$, System
@@ 15                   $\mathcal{M}_{n,k}\leftarrow UpdateMemory(\mathcal{M}_{n,k-1},$ $u_{n,k},d_{n},t)$;
17                         $a_{n,m}\leftarrow SystemAns(\mathcal{M}_{n,m},$ $q_{n,m,c},d_{n},t)$;
@@ 4.1 Simulator
Algorithm 1 outlines the simulation process of DialSim, designed to emulate a real-time conversation. In this simulator, each participant’s utterance (including the system’s) occurs at a predefined time interval (same as time limit), and the system should update its memory within this interval. [^5] The memory can be incrementally updated in various ways (e.g., by storing each utterance separately or by summarizing the session up to the current utterance). A detailed discussion of these methods is provided in § 5.2. If updating the memory is not completed within the interval, the simulator will move on to the next utterance (Line 13). During the simulation, other characters ask questions (selected from LongDialQA) to the system (Line 8-10), except in sessions where the system is the only one talking (Line 5-6). The timing to ask a question is chosen randomly within the session (Line 8), and the speaker who asks the question is also chosen randomly. However, to make the simulation realistic, it is crucial to ensure that the chosen speaker is still present and hasn’t left the session. We achieved this by randomly choosing from characters who were present within three turns of the system’s last utterance (Line 9). Then, a question is randomly selected and asked in the style of the corresponding speaker (Line 10). The system then must respond to the question using its memory, all within the time limit (Line 15). The prompt for the response is created by combining the question with the dialogue history stored in the memory. If the response is not completed within the time limit, it will be considered a failure, and the simulator will move on to the next utterance. The prompt we used is provided in Appendix F.
@@ 5 Experiments
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 38.53 (0.89) 32.65 (2.65) 49.04 (1.67) 40.27 (1.36) 40.10 (0.75) 44.36 (2.36) 42.53 (1.26) GPT-3.5 31.82 (1.31) 25.58 (1.78) 39.70 (1.86) 32.09 (0.84) 32.06 (1.60) 36.84 (1.77) 36.69 (1.25) Gemini 1.0 pro 2.96 (0.31) 28.77 (1.83) 25.07 (2.40) 35.27 (1.80) 34.22 (0.49) 31.83 (0.41) 35.75 (2.93) Open Tülu 2-70B 0.37 (0.15) 20.94 (0.75) 20.27 (0.99) 19.75 (0.08) 31.76 (1.84) 10.15 (0.55) 18.87 (0.30) Tülu 2-7B 0.84 (0.15) 12.68 (0.24) 19.58 (1.04) 26.84 (0.85) 14.08 (0.89) 17.39 (1.37) 25.21 (1.28) Llama3.1-70B 0.60 (0.06) 31.08 (1.21) 0.55 (0.12) 16.26 (5.05) 39.00 (0.30) 2.26 (0.42) 20.14 (0.22) Llama3.1-8B 28.82 (1.94) 27.12 (0.95) 34.14 (0.85) 30.91 (0.63) 29.76 (1.31) 33.25 (0.57) 24.48 (0.60) Mixtral-8x7B 1.88 (0.26) 16.84 (0.95) 26.23 (0.90) 17.11 (1.94) 17.94 (1.32) 26.78 (1.04) 15.40 (1.39) Mistral-7B 2.82 (0.46) 24.22 (2.04) 33.07 (1.01) 29.29 (1.76) 28.30 (1.93) 29.15 (1.67) 25.41 (1.53) Gemma-7B 16.60 (0.84) 22.11 (1.73) 24.30 (2.04) 18.33 (1.37) 26.42 (2.48) 22.54 (0.78) 18.80 (0.64) Gemma-2B 0.68 (0.20) 24.06 (2.03) 24.22 (1.34) 25.79 (1.00) 25.31 (1.55) 24.48 (1.62) 25.78 (1.12) Random Guessing 20.00 20.00 20.00 20.00 20.00 20.00 20.00
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@ 5.1 Experimental Setting
We provide the option to choose between multiple-choice and open-ended question formats. In our experiments, we used the multiple-choice format to efficiently and accurately assess the system’s dialogue understanding capabilities. Table 1 presents the statistics for LongDialQA, highlighting a significant disparity between the number of answerable and unanswerable questions. To ensure a balanced distribution of correct answers during the simulation, 20% of the questions were intentionally designed to be unanswerable, with each question providing five possible choices.
@@ 5.2 Baselines
We experimented with representative LLM-based conversation systems capable of handling long-term dialogue, focusing on two memory management approaches. The first method, namely Base LLM, simply prepends as many of the most recent utterances as the model’s context length allows. The second method, namely RAG-based, employs a retriever to search for relevant dialogue history from the system’s memory (external storage) and includes it in the prompt (Lewis et al., 2020). This method can be broken down into three ways for storing dialogue history: each speaker’s utterance individually, the entire session, and a summarized version of each session (denoted as Utterance, Session Entire, and Session Sum. in Table 2). The retrieval from the memory was performed using BM25 (Robertson et al., 2009) and cosine similarity with the OpenAI embeddings (OpenAI, 2024c).
For the LLMs, we used both API-based models (i.e., Gemini-1.0 Pro, 1.5 Pro (Team et al., 2023; Reid et al., 2024), Claude 3 Opus (Anthropic, 2024), GPT-3.5, 4o, 4o-mini (OpenAI, 2023b, 2024b, 2024a)) and open-source models (i.e., Tülu 2-7B, 70B (Ivison et al., 2023), Llama3.1-8B, 70B (Meta, 2024), Mistral-7B, 8x7B (Jiang et al., 2023, 2024), and Gemma-2B, 7B (Team et al., 2024)). [^6] Gemini-1.5 Pro, Claude 3 Opus, and GPT-4o were evaluated only in the BM25-Session Entire and oracle setting to measure their performance upper bound due to their high prices. The experimental results can be found in Appendix K. To emulate conversational settings, we used chat templates for instruction-tuned models or directly used chat models.
@@ 5.3 Results
Overall Performance   Table 2 shows that API-based models outperformed open-source models due to their superior inference capabilities and faster response times in our setting. However, the performances of all baselines were below 50%, suggesting that current LLMs have limitations in their ability to serve as conversation systems for long-term multi-party dialogues. The experimental results for Friends, The Big Bang Theory, and The Office exhibited similar trends. The detailed results are described in Appendix H.
Model Base LLM BM25 OpenAI Embedding UT SE SS UT SE SS GPT-4o-mini 38.91 34.44 49.21 42.23 38.91 43.64 42.40 GPT-3.5 31.81 26.91 39.45 32.77 32.41 35.78 35.98 Gemini 1.0 pro 28.36 28.10 39.90 34.11 34.26 30.93 33.96 Llama3.1-70B 36.36 31.84 43.17 43.81 39.85 43.17 48.49 Llama3.1-8B 28.78 29.89 34.70 33.93 31.63 32.91 35.59 Mixtral-8x7B 42.19 31.84 46.47 32.31 35.51 41.24 34.18 Mistral-7B 32.93 28.20 35.09 30.16 30.12 31.00 30.10 $\dagger$: Limited the maximum context length to 8k tokens.
@@
Storing the entire session consistently outperforms other history storing methods, as shown in Table 3. This is because individual utterances lack adequate context, and crucial information may be lost during summarization. However, Llama3.1 models achieved the best performance when using Session Sum. as a history saving method, owing to their strong summarization capabilities. Additionally, contrary to our expectations, Mixtral’s Base LLM (i.e., without history retrieval) outperforms some retrieval-based models in settings with unlimited time. This is due to Mixtral’s context length of 32k tokens, which is long enough to accommodate half a season of the script, allowing it to utilize a longer dialogue history than some of the other baselines. However, in a setting with a time limit, Mixtral’s performance significantly drops due to its long inference time. Therefore, for a conversation system to converse in real-time, it is necessary to select a reasonably appropriate length of dialogue history.
Advanced techniques for storing and retrieving history are essential to engage in long-term multi-party dialogues. We conducted experiments under the oracle setting, where systems were given evidence sessions along with their dates (see Figure 2). Under these conditions, Llama3.1-70B achieved a top performance of 69.86% in an unlimited time scenario, outperforming the best RAG-based method by 21.37%. This significant performance gap highlights the importance of effective memory management techniques. Detailed experimental results are provided in Appendix K.
TKG-based questions present a greater challenge than fan quiz-based ones, with two-hop questions being particularly difficult. To assess the difficulty levels across different question types, we conducted an error analysis on GPT-4o-mini, based on BM25-Session Entire, which showed the highest performance. The results showed that fan quiz-based questions had an accuracy of 58.80%, while TKG-based questions scored lower at 46.40%, highlighting the greater difficulty of TKG-based questions. Breaking down TKG-based questions further, one-hop questions had a performance of 66.67%, whereas two-hop questions had a performance of 13.53%, underscoring the challenge of two-hop questions. Furthermore, even in the oracle setting, while the performance of one-hop questions increased to 84.05%, two-hop questions remained at 28.45%. This indicates that two-hop questions are challenging not only in terms of history retrieval but also in reasoning across the given sessions.
Model Original Swapping Names New Names GPT-4o-mini 49.21 38.40 ($\downarrow$10.81) 46.57 ($\downarrow$2.64) Llama3.1-70B 48.49 39.00 ($\downarrow$9.49) 44.27 ($\downarrow$4.22) Llama3.1-8B 35.59 31.59 ($\downarrow$4.00) 32.23 ($\downarrow$3.36) Mixtral-8x7B 46.47 37.72 ($\downarrow$8.75) 39.98 ($\downarrow$6.49) Mistral-7B 35.09 30.65 ($\downarrow$4.44) 34.59 ($\downarrow$0.50)
@@
The adversarial versions of LongDialQA are effective for the rigorous dialogue understanding evaluation of conversation systems. We conducted further experiments using DialSim on the adversarial versions of LongDialQA. The results shown in Table 4 indicated a non-trivial drop in overall performance compared to the original setup. This decline can be attributed to the systems relying not only on the dialogue history but also likely on their pre-trained knowledge when answering questions. Additionally, the performance decrease was more pronounced when names were swapped compared to when new names were assigned. This suggests that new names represent new information, while swapped names in the dialogue history conflicted with the pre-trained knowledge, leading to reduced reasoning performance. Detailed experimental results are provided in Appendix L.
@@ 6 Conclusion
In this paper, we introduce DialSim, a simulator designed for evaluating the ability of conversation systems to understand long-term, multi-party dialogues in real-time settings. To run DialSim, we first constructed LongDialQA, a dataset based on dialogues from well-known TV show scripts. LongDialQA also includes questions derived from fan quizzes and a temporal knowledge graph, enabling a comprehensive assessment of conversation systems. Using DialSim, we evaluated the latest conversation systems and uncovered significant limitations in their ability to effectively handle complex, multi-party, long-term dialogues in real-time scenarios.
@@
Despite its strengths, our simulator has two main limitations. First, while the questions and answers are logically paired for accurate evaluation, the random selection of questions could introduce a bit of awkwardness during conversations. Second, while we considered incorporating industry-specific dialogues such as chat logs from customer service or retail, where conversation systems could be used for business purposes, these dialogue datasets are usually proprietary and not publicly accessible. In future developments, we will focus on enhancing the natural flow of interactions and creating simulators that are applicable to real-world industries.
@@ Appendix A LLM’s Prior Knowledge of the TV shows
We asked GPT-4o to explain the plot of specific episodes of Friends. It accurately described the plots, as shown in Figure 45. Notably, it provided these answers without any web browsing, suggesting that GPT-4o might have learned about these TV shows during its pre-training process.
@@ Appendix C Question Generation Based on Fan Quizzes
First, we collected quizzes for each season and episode of Friends, The Big Bang Theory, and The Office from the FunTrivia website. For each episode $p_{i}$ in $Script_{pre}$, we used GPT-4 to determine if the crawled questions ${CrQ}_{i}=\{q_{i,0},q_{i,1},...,q_{i,l}\}$ could be answered using only $p_{i}$. If a question $q_{i,m}$ could be answered, GPT-4 identified the scenes $ES_{i,m}$ that provide evidence for the answer, compiling them into $Q_{i}=\{(q_{i,m},ES_{i,m})\}_{m=0}^{l}$. Subsequently, the authors reviewed each $ES_{i,m}$, made necessary corrections, and annotated whether a single scene from $ES_{i,m}$ was sufficient to answer $q_{i,m}$ or if multiple scenes were needed to be considered simultaneously. For each $s_{i,k}$ within $p_{i}$, we assessed the answerability of the questions in $Q_{i}$.